OpenStudy (turtwig):

I'm not sure how to solve this equation 3(2y−1)−y=2(y+2)−(y−4) I'm trying to determine whether there is one solution, infinite solutions, or no solutions. The parenthesis are confusing me a little, could somebody give me step-by-step instructions?

OpenStudy (avi13):

3(2y−1)−y=2(y+2)−(y−4) == 6y-3-y=2y+4-y+4 5y-3=y+8 4y=11 y=11/4

OpenStudy (mecharv):

Since no two of the y terms are being multiplied you have only a linear equation with one variable. First let is solve the Right Hand side. \[3(2y-1)-y=6y-3-y=5y-3\] Now the left hand side \[2(y+2)-(y-4)=2y+4-y+4=y+8\] Equating LHS and RHS, \[y+8=5y-3\] \[y-5y=-3-8\] \[-4y=-11\] \[y=\frac{ -11 }{ -4 }\] \[y=\frac{ 11 }{ 4 }\] SO clearly you have only one solution. Cheers :)

OpenStudy (mecharv):

@turtwig Still confused?

OpenStudy (turtwig):

@mecharv I think I understand now :D Thank you for the thorough response!

OpenStudy (mecharv):

Anytime :)